# Exploiting Observation Bias to Improve Matrix Completion

### v2

@@ 2.2 Rationale behind the proposed model
The model we propose captures a phenomenon that often occurs in practice. As an example, consider the case of the Glance smart lock-screens, where the rows of $\mathbf{X}$ correspond to users, the columns correspond to individual glances, and an entry $x_{ij}$ corresponds to the time user $i$ spends interacting with glance $j$. We note here that a user not interested in sports is unlikely to spend time viewing a glance recounting a recent basketball game. Moreover, based on the behavior of this user, the recommendation algorithm is unlikely to recommend glances about sports. Hence, information about the users’ preferences can be extracted both from their engagement time and which lock screens they are shown. This shared information corresponds to common latent features that affect both the observation pattern and the outcome in our model.
Low-rank representation.
In view of the above discussion, we may hypothesize that both the observation probabilities and outcomes of interest may be expressed as $p_{ij}=g(\mathbf{z}_{i},\mathbf{w}_{j})$ and $x_{ij}=h(\mathbf{z}_{i},\mathbf{w}_{j})$ for all $i,j\in[n]$, where $g$ and $h$ are both “nice" functions of shared latent factors $\mathbf{z}_{i},\mathbf{w}_{j}$ in $\mathbb{R}^{d}$. Here, by “nice" functions, we could mean for instance that $g$ and $h$ are bounded and Lipschitz. This is a reasonable assumption under which we may use a classical spectral decomposition Conway (2019). Below, we present a variant of this decomposition which we borrow from Shah et al. (2020) (see their Theorem 1).
Theorem 2.1.
@@
Let $h:\mathcal{S}^{d-1}\times\mathcal{S}^{d-1}\mapsto\mathbb{R}$ be bounded and $L$-Lipschitz with respect to the $1$-product metric in $\mathcal{S}^{d-1}\times\mathcal{S}^{d-1}$. Then, there exists a non-increasing sequence $(\sigma_{k})_{k\geq 1}$ with $\sum_{k=1}^{\infty}\sigma_{k}^{2}<\infty$, and orthogonal families of functions $\{\alpha_{k}\in L^{2}(\mathcal{S}^{d-1}):k\in\mathbb{N}^{\star}\}$, and $\{\gamma_{k}\in L^{2}(\mathcal{S}^{d-1}):k\in\mathbb{N}^{\star}\}$ such that, for all $\mathbf{z},\mathbf{w}\in\mathcal{S}^{d-1}$,
$h(\mathbf{z},\mathbf{w})=\sum_{k=1}^{\infty}\sigma_{k}\alpha_{k}(\mathbf{z})\gamma_{k}(\mathbf{w}).$ (4)
@@
Now using Theorem 2.1, we may argue that both $g$ and $h$ admit a low-rank representation, say of rank $d$ for simplicity, where we write $p_{ij}=g(\mathbf{z}_{i},\mathbf{w}_{j})=\sum_{\ell=1}^{d}u_{\ell}(\mathbf{z}_{i})v_{\ell}(\mathbf{w}_{j})$, and $x_{ij}=h(\mathbf{z}_{i},\mathbf{w}_{j})=\sum_{\ell=1}^{d}\alpha_{\ell}(\mathbf{z}_{i})\gamma_{\ell}(\mathbf{w}_{j})$ for all $i,j\in[n]$ with $u(\cdot),v(\cdot),\alpha(\cdot)$, and $\gamma(\cdot)$ are mappings from $\mathcal{S}^{d-1}$ to $\mathbb{R}^{d}$. Thus, if additionally $u(\cdot)$ and $v(\cdot)$ are invertible, then we may define $\theta=\alpha\circ u^{-1}$ and $\phi=\gamma\circ v^{-1}$ and obtain $p_{ij}\propto\mathbf{u}_{i}^{\top}\mathbf{v}_{j}$ and $x_{i,j}=f(\mathbf{u}_{i},\mathbf{v}_{j})=\sum_{\ell=1}^{d}\theta_{\ell}(\mathbf{u}_{i})\phi_{\ell}(\mathbf{v}_{j})$, where $\mathbf{u}_{i}=u(\mathbf{z}_{i})$ and $\mathbf{v}_{j}=v(\mathbf{w}_{j})$ for all $i,j\in[n]$. With this rough argument, we recover a low-rank representation similar to the one described in our model.
3 The Mask Nearest Neighbor (MNN) Algorithm
We propose the Mask Nearest Neighbor (MNN) algorithm that runs in two stages: (i) The first stage consists in reducing the noise in the observed outcomes by clustering users and items that are sufficiently close to each other, then averaging over these clusters whenever there are enough entries observed within a cluster. This averaging leads to a new matrix with a reduced size and still missing entries; (ii) The second stage consists in imputing the missing entries of the reduced matrix, then extrapolating from it to predict all outcomes.
3.1 Clustering & Denoising
Distance estimation.
@@ 3.1 Clustering & Denoising
The first step of the algorithm is to estimate the user-user and item-item distances in the latent space, denoted $d_{ij}^{\mathbf{u}}=\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}$, and $d_{ij}^{\mathbf{v}}=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|_{2}$, for all $i,j\in[n]$, respectively. To that end, we start by constructing the centered matrix of observations $\widetilde{\mathbf{A}}=\mathbf{A}-(\rho_{n}/2)\bf 1\bf 1^{\top}$, which we then project onto the space of rank $d$ matrices. More precisely, given the singular value decomposition $\widetilde{\mathbf{A}}=\sum_{i=1}^{n}\sigma_{i}\mathbf{q}_{i}\mathbf{w}_{i}^{\top}$ we construct an estimate of $\widetilde{\mathbf{P}}=\mathbb{E}[\widetilde{\mathbf{A}}]$ as follows:
@@
$\hat{d}^{\mathbf{u}}_{ij}$ $=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}\left\lVert\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widehat{\widetilde{\mathbf{P}}}_{j,:}\right\rVert_{2},$
@@
Clustering.
We group the users and items by first selecting maximal subsets that are sufficiently separated. Concretely, we initialize the central users, denoted $\mathcal{U}^{\star}\subseteq[n]$, as the empty set. Then, we repeatedly add an arbitrary user to the set if its estimated distance from any other central user is at least $6\varepsilon_{n}$, where $\varepsilon_{n}$ is a hyperparameter we defer quantifying. $\mathcal{V}^{\star}$ is defined analogously.
By assigning each non-central user to its closest central counterpart in estimated distance, a partition of the users $\bigcup_{i\in\mathcal{U}^{\star}}\mathcal{U}_{i}=[n]$ is formed. Similarly construct the item partition $\bigcup_{j\in\mathcal{V}^{\star}}\mathcal{V}_{j}$. Note that we index the partitions by the sets $\mathcal{U}^{\star},\mathcal{V}^{\star}$ to avoid confusion from multiple indexing systems.
@@
Denoising.
At this stage, we are ready to define the ground truth clustered outcome matrix $\mathbf{H}=(h_{ij})_{i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}}$, where
$h_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j}),\quad i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}.$ (7)
We again emphasize that the rows and columns of $\mathbf{H}$ are indexed by central users $\mathcal{U}^{\star}$ and central items $\mathcal{V}^{\star}$ for simplicity. Using data, we form the unfilled clustered outcome matrix $\overline{\mathbf{H}}\in\mathbb{R}^{|\mathcal{U}^{\star}|\times|\mathcal{V}^{\star}|}$, where for all $(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$,
@@ 3.2 Imputation & Outcome Prediction
We will impute the missing entries of $\overline{\mathbf{H}}$ using the available entries. We will define the imputed clustered outcome matrix $\widehat{\mathbf{H}}$, with
$\hat{h}_{ij}=\begin{cases}\bar{h}_{ij}&\quad\text{if $\bar{h}_{ij}\neq\star$};\\ \texttt{Impute}(\overline{\mathbf{H}},i,j)&\quad\text{otherwise.}\end{cases}$ (9)
The subroutine $\texttt{Impute}(\overline{\mathbf{H}},i,j)$ proceeds as follows:
@@
    First, we search over the sets $\mathcal{U}^{\star}$ and $\mathcal{V}^{\star}$ to find subsets of $d$ distinct central users $\mathcal{I}$ and $d$ distinct central items $\mathcal{J}$ such that $\sigma_{r}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})$ is maximized and
    $\!\!\!\!\!\!\!\!\!\!\!\!\!\!\forall(i^{\prime},j^{\prime})\in\mathcal{I}\times\mathcal{J},\quad\bar{h}_{i,j^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime},j^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime},j}\neq\star.$ (10)
@@
    $\!\!\!\!\!\!\!\!\!\!\!\!\texttt{Impute}(\overline{\mathbf{H}},i,j)\!=\!\!\begin{cases}\!\overline{\mathbf{H}}_{i,\mathcal{J}}\!\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\!\right)^{\!\dagger}\!\overline{\mathbf{H}}_{\mathcal{I},j}&\!\!\!\!\text{if }\mathcal{I},\mathcal{J}\text{ are found};\\ \star&\!\!\!\!\text{otherwise.}\end{cases}$ (11)
@@
Once $\widehat{\mathbf{H}}$ is constructed, predicting outcomes is simple: For any $i,j\in[n]$, there exists $k\in\mathcal{U}^{\star}$ and $\ell\in\mathcal{V}^{\star}$ such that $i\in\mathcal{U}_{k}$ and $j\in\mathcal{V}_{\ell}$, we then set $\hat{x}_{ij}=\hat{h}_{kl}$.
We have now completed the description of MNN. For the reader’s convenience, we provide a pseudocode of our algorithm in Appendix G.
4 Prediction Error Guarantee
We analyze the performance of MNN in terms of its prediction error at a user-item level. To that end, our analysis relies on three key ingredients: First, we show in that the our joint clustering &\&& denoising procedure yields good estimates for the outcomes of the central user-item pairs whenever they are revealed (see Proposition 4.1); Second, we characterize the pattern of the revealed central user-item pairs (see Proposition 4.2); Third, we show that the imputed central user-item pairs are accurate (4.3). In conclusion, we provide a finite-sample guarantee on the prediction error of MNN in Theorem 4.1, and discuss the obtained rates in comparison with the optimal parametric rates for supervised learning parametric.
4.1 Analysis
Clustering &\&& Denoising.
Let us denote the set of revealed user-item clusters as follows
$\Omega^{\star}=\{(i,j):i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star},\quad|\Omega_{ij}|\geq N_{n}\},$ (12)
In Proposition 4.1, we make the choice of $N_{n}$ precise and provide a finite-sample guarantee on the error rate between the revealed entries of $\overline{\mathbf{H}}$ and $\mathbf{H}$.
Proposition 4.1.
@@ 4.1 Analysis
holds with probability at least $1-\delta$, provided $n\gtrsim d\log\left(e\,n/\delta\right)$ and $\varepsilon_{n}$ satisfies
@@
The precise statement of Proposition 4.1 is given in Lemma C.1, and its proof in Appendix C.
Observations Pattern of $\overline{\mathbf{H}}$.
In order to ensure correctness of the subroutine $\texttt{Impute}(\overline{\mathbf{H}},i,j)$, the mappings $\theta(\cdot)$ and $\phi(\cdot)$ must be “non-degenerate” in some appropriate sense so that the search for $\mathcal{I}$ and $\mathcal{J}$ is successful, and that the smallest singular value of $\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}$ is bounded away from 00 independently from $n$. This will in turn ensure that observation bias may still be exploited to impute the missing entries. Below, we precisely state what we mean by “non-degenerate”.
Definition 4.1 (Non-degeneracy).
@@
Now, we are ready to present Proposition 4.2, in which we describe the pattern of revealed entries in $\overline{\mathbf{H}}$ provided that $\theta(\cdot)$ and $\phi(\cdot)$ are “non-degenerate”. This in turn ensures that $\texttt{Impute}(\overline{\mathbf{H}},i,j)$ will successfully terminate with high probability.
Proposition 4.2.
Assume that $\theta(\cdot)$ and ϕ()italic-$\phi(\cdot)$ are non-degenerate with parameters $m,\kappa>0$, and that $N_{n}$ is chosen as in Proposition 4.1. Let $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ be such that $\bar{h}_{i,j}=\star$, (i.e., $(i,j)$ is a non-revealed user-item cluster in $\overline{\mathbf{H}}$). Then, for all $\delta\in(0,1)$, the event: there exists $\mathcal{I}\subset\mathcal{U}^{\star},\mathcal{J}\subseteq\mathcal{V}^{\star}$ with $|\mathcal{I}|=|\mathcal{J}|=d$, such that
[…]
@@
The proof of Proposition 4.3 is given in Appendix D. The precise statement of the proposition is given in D.1. It is worth noting that the “non-degeneracy” property allows us to rule out pathological cases for which it is impossible to impute the missing entries from the observed one, at least using our approach. The need for such property is apparent in the proof, and we refer the reader to Appendix E to see how this property comes into our analysis.
Matrix completion.

### v3

@@ 2.2 Why this model makes sense?
The motivation behind the model described in §2.1 is precisely to capture a phenomenon that often occurs in practice, namely that of confounding. An example of this is the case of the Glance smart lock-screens, where the rows of $\mathbf{X}$ correspond to users, the columns correspond to individual glances and an entry $x_{ij}$ corresponds to the time user $i$ spends interacting with glance $j$. We note here that a user not interested in sports is unlikely to spend time viewing a glance recounting a recent basketball game. Moreover, based on the behavior of this user, the recommendation algorithm is unlikely to recommend glances about sports. Hence, information about the users’ preferences can be extracted both from their engagement time and which lock screens they are shown. This shared information corresponds to common latent features that affect both the observation pattern and the outcome in our model.
A generic confounding-aware MNAR framework.
Formally, we may consider that the common features as discussed above for a user-item pair $(i,j)$ may be denoted by the two $d$-dimensional vectors $\mu_{i}$ and $\nu_{j}$. Then, as posited before, $\mu_{i}$ and $\nu_{j}$ influence both the the probability $p_{ij}$, corresponding to observing the outcome of interest for the user-item pair $(i,j)$, as well as the outcome of interest itself $x_{ij}$:
$p_{ij}$ $=g(\mu_{i},\nu_{j}),$ (4)
$x_{ij}$ $=h(\mu_{i},\nu_{j}),$ (5)
where $g$ (resp. $h$) is a function that maps the common factors to a value in $[0,1]$ (resp. in some bounded domain of possible outcome values). Without further assumptions on $g$ and $h$, the task of matrix completion may be unsolvable. It therefore natural to assume such functions to be nice in some appropriate sense. The model we described here is somewhat general and is of independent interest. We refer to a concurrent work [22] where a similar shared latent factored model was considered in the context of transfer learning between biological networks.
The low-rank hypothesis.
It is reasonable to consider that $g$ and $h$ are nice functions in the sense that they are bounded and Lipschitz. In this case, we may use a classical functional decomposition result [14]. We present below a variant of such decomposition (see Theorem 1 in [37]):
Proposition 1.
@@
Let $g:\mathcal{S}^{d-1}\times\mathcal{S}^{d-1}\mapsto\mathbb{R}$ be bounded and $L$-Lipschitz with respect to the $1$-product metric in $\mathcal{S}^{d-1}\times\mathcal{S}^{d-1}$. Then, there exists a non-increasing sequence $(\sigma_{k})_{k\geq 1}$ with $\sum_{k=1}^{\infty}\sigma_{k}^{2}<\infty$, and orthogonal families of functions $\{\alpha_{k}\in L^{2}(\mathcal{S}^{d-1}):k\in\mathbb{N}^{\star}\}$, and $\{\gamma_{k}\in L^{2}(\mathcal{S}^{d-1}):k\in\mathbb{N}^{\star}\}$ such that, for all $\mathbf{z},\mathbf{w}\in\mathcal{S}^{d-1}$,
$g(\mathbf{z},\mathbf{w})=\sum_{k=1}^{\infty}\sigma_{k}\alpha_{k}(\mathbf{z})\gamma_{k}(\mathbf{w}).$ (6)
@@
Using Proposition 1, it is reasonable to assume that both $g$ and $h$ admit a low-rank representation, say of rank $d$ for simplicity, where we write $p_{ij}=g(\mu_{i},\nu_{j})\approx\sum_{\ell=1}^{d}a_{\ell}(\mu_{i})b_{\ell}(\nu_{j})$ and $x_{ij}=h(\mu_{i},\nu_{j})\approx\sum_{\ell=1}^{d}\alpha_{\ell}(\mu_{i})\beta_{\ell}(\nu_{j}),$ where $\alpha(\cdot),\beta(\cdot),a(\cdot),b(\cdot)$ are mappings from $\mathcal{S}^{d-1}$ to $\mathbb{R}^{d}$. Thus, if additionally $a(\cdot)$ and $b(\cdot)$ are invertible, then we may define $\theta=\alpha\circ a^{-1}$ and $\phi=\beta\circ b^{-1}$, and finally obtain (with appropriately defined $\mathbf{u}_{i},\mathbf{v}_{j}$ with $i,j\in[n]$)
$p_{ij}$ $\approx\mathbf{u}_{i}^{\top}\mathbf{v}_{j},$ (7)
$x_{ij}$ $\approx\sum_{\ell=1}^{d}\theta_{\ell}(\mathbf{u}_{i})\phi_{\ell}(\mathbf{v}_{j}).$ (8)
We clearly see the similarity between (1) (resp. (2)) and (8) (resp. (7)). Therefore, we see the model we described in §2.1 as a natural and fairly general instance of a an MNAR model that accounts for the presence of confounding.
3 Mask Nearest Neighbor (MNN)
In this section, we present the Mask Nearest Neighbors (MNN) algorithm, which gets its name from its two stages: (i) Learning Distances From Masks: estimate distances between the latent factors associated with users, items using mask matrix $\mathbf{A}$, and using them to cluster users, items; (ii) Estimation using Nearest Neighbor: de-noising observed values and impute missing values in the outcome matrix via non-parametric nearest neighbor supervised learning.
3.1 Distance Estimation, Clustering
@@
Distance estimation. The first step of the algorithm is to estimate the user-user and item-item distances in the latent space, denoted $d_{ij}^{\mathbf{u}}=\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}$, and $d_{ij}^{\mathbf{v}}=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|_{2}$, for all $i,j\in[n]$, respectively. To that end, we start by constructing the centered matrix of observations $\widetilde{\mathbf{A}}=\mathbf{A}-(\rho_{n}/2)\bf 1\bf 1^{\top}$, which we then project onto the space of rank $d$ matrices. More precisely, given the singular value decomposition $\widetilde{\mathbf{A}}=\sum_{i=1}^{n}\sigma_{i}\mathbf{q}_{i}\mathbf{w}_{i}^{\top}$ we construct an estimate of $\widetilde{\mathbf{P}}=\mathbb{E}[\widetilde{\mathbf{A}}]$ as follows:
@@
$\hat{d}^{\mathbf{u}}_{ij}$ $=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}\left\lVert\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widehat{\widetilde{\mathbf{P}}}_{j,:}\right\rVert_{2},$ (10)
@@
In above, $\widetilde{\mathbf{P}}_{i,:}$ (resp. $\widetilde{\mathbf{P}}_{:,i}$) correspond to $i$th row (resp. $i$th column) of $\widetilde{\mathbf{P}}$.
Clustering. The goal is to cluster users (resp. items) using their distance estimates so that any two users (resp. items) that belong to the same cluster are within distance $\varepsilon_{n}>0$, a hyper-parameter of choice. Towards that, first we identify central users, denoted as $\mathcal{U}^{\star}\subset[n]$, that are farther apart from each other. To identify $\mathcal{U}^{\star}$, start with it to be empty set. We repeatedly add an arbitrary user to the set if its estimated distance from all user in $\mathcal{U}^{\star}$ is at least $6\varepsilon_{n}$. Stop when no more users can be added. Similarly construct $\mathcal{V}^{\star}$ for items.
We shall create cluster of users around each central user in $\mathcal{U}^{\star}$. To that end, for each $i\in[n]\backslash\mathcal{U}^{\star}$, associate it to the closest of the users in $\mathcal{U}^{\star}$ as per the distance estimated. Thus, each $i\in\mathcal{U}^{\star}$ corresponds to a central user, and let $\mathcal{U}_{i}\subset[n]$ be the users associated with it. Thus we have a partition or clusters of the users $\bigcup_{i\in\mathcal{U}^{\star}}\mathcal{U}_{i}=[n]$. Similarly construct the item partition $\bigcup_{j\in\mathcal{V}^{\star}}\mathcal{V}_{j}$. Note that we abuse notation here by indexing the partitions by the elements of $\mathcal{U}^{\star},\mathcal{V}^{\star}$ to avoid overload of notation.
@@
Figure 1: Visual depiction of the first stage of MNN. The matrix of outcomes $\mathbf{Y}$ is illustrated by figure (a). From the outcomes matrix $\mathbf{Y}$, we extract the matrix $\mathbf{A}$ which is fully observed and illustrated by figure (b). The matrix $\mathbf{A}$ is then used to learn distances between the shared latent user factors (resp. latent item factors), then use these distances to cluster users (resp. items). Using these clusters, we may rearrange the rows and columns of the outcome matrix as illustrated by figure (c), where each block correspond to the Cartesian product of a cluster of users and a cluster of items.
3.2 Outcome Prediction
Denoising. Define the ground truth clustered outcome matrix $\mathbf{H}=(h_{ij})_{i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}}$, where $h_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j}),\quad i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}.$ We again emphasize that the rows and columns of $\mathbf{H}$ are indexed by central users $\mathcal{U}^{\star}$ and central items $\mathcal{V}^{\star}$ for simplicity. Using data, we form the unfilled clustered outcome matrix $\overline{\mathbf{H}}\in\mathbb{R}^{|\mathcal{U}^{\star}|\times|\mathcal{V}^{\star}|}$, where for all $(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$,
@@ 3.2 Outcome Prediction
We will impute the missing entries of $\overline{\mathbf{H}}$ using the available ones. The resulting matrix of this imputation, also referred to as imputed clustered outcome matrix, will be denoted by $\widehat{\mathbf{H}}$. Let $i,j\in[n]$, if $\bar{h}_{ij}\neq\star$, then we simply set $\hat{h}_{ij}=\bar{h}_{ij}$. If $\bar{h}_{ij}=\star$, then we impute the missing entry by proceeding as follows:
@@
    First, we search over the sets $\mathcal{U}^{\star}$ and $\mathcal{V}^{\star}$ to find subsets of $d$ distinct central users $\mathcal{I}$ and $d$ distinct central items $\mathcal{J}$ such that $\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})$ is maximized and
    $\forall(i^{\prime},j^{\prime})\in\mathcal{I}\times\mathcal{J},\ \ \bar{h}_{i,j^{\prime}}\neq\star,\bar{h}_{i^{\prime},j^{\prime}}\neq\star,\bar{h}_{i^{\prime},j}\neq\star.$ (13)
@@
    $\hat{h}_{ij}=\begin{cases}\overline{\mathbf{H}}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\overline{\mathbf{H}}_{\mathcal{I},j}&\text{if }\mathcal{I},\mathcal{J}\text{ are found};\\ \star&\text{otherwise.}\end{cases}$ (14)
As we shall show later, $(\mathcal{I},\mathcal{J})$ are guaranteed to be found, thus ensuring the correctness of the imputation procedure.
Figure 2: Visual depiction of the second stage of MNN. The blue cells correspond to observed or imputed entries, and the red cells correspond to unobserved entries. The matrix illustrated in (a) corresponds to the outcomes matrix after rearranging its rows and columns. The matrix in (b) corresponds to the resulting matrix after the denoising step using (12). The matrix illustrated in (c) corresponds to the filled entries after using our the imputation procedure using (14).
@@
Once $\widehat{\mathbf{H}}$ is constructed, predicting outcomes is simple: For any $i,j\in[n]$, there exists $k\in\mathcal{U}^{\star}$ and $\ell\in\mathcal{V}^{\star}$ such that $i\in\mathcal{U}_{k}$ and $j\in\mathcal{V}_{\ell}$, we then set $\hat{x}_{ij}=\hat{h}_{kl}$. This completes the description of MNN.
4 Prediction Guarantee of MNN
Our main result is Theorem 1 which corresponds to the finite-sample performance guarantee of MNN in terms of its prediction error at the user-item level. To establish this result, our analysis relies on three key ingredients: First, we show in that the joint clustering &\&& denoising procedure yields good estimates for the outcomes of the central user-item pairs whenever they are revealed (see Proposition 2); second, we characterize the pattern of the revealed central user-item pairs (see Proposition 3); third, we show that the imputed central user-item pairs are accurate (Proposition 4).
4.1 Clustering, Denoising.
We denote the set of revealed user-item clusters by $\Omega^{\star}=\{(i,j):i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star},\quad|\Omega_{ij}|\geq N_{n}\}.$ In Proposition 2, we make the choice of $N_{n}$ precise and provide a finite-sample guarantee on the error rate between the revealed entries of $\overline{\mathbf{H}}$ and $\mathbf{H}$.
Proposition 2.
@@ 4 Prediction Guarantee of MNN
holds with probability at least $1-\delta$, provided $n\gtrsim d\log\left(e\,n/\delta\right)$ and
@@
We prove Proposition 2 in Appendix A.
4.2 Observations Pattern of $\overline{\mathbf{H}}$.
In order to ensure correctness of the procedure by which we construct $\widehat{\mathbf{H}}$, the mappings $\theta(\cdot)$ and $\phi(\cdot)$ must be “non-degenerate” in some appropriate sense so that the search for $\mathcal{I}$ and $\mathcal{J}$ is successful, and that the smallest singular value of $\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}$ is bounded away from 00 independently from $n$. This will in turn ensure that observation bias may still be exploited to impute the missing entries. Below, we precisely state what we mean by “non-degenerate”.
Definition 1 (Non-degeneracy).
@@
Now, we are ready to present Proposition 3, in which we describe the pattern of revealed entries in $\overline{\mathbf{H}}$, provided that $\theta(\cdot)$ and $\phi(\cdot)$ are “non-degenerate”. This in turn ensures that $\texttt{Impute}(\overline{\mathbf{H}},i,j)$ will successfully terminate with high probability.
Proposition 3.
Assume that $\theta(\cdot)$ and $\phi(\cdot)$ are non-degenerate with parameters $m,\kappa>0$, and that $N_{n}$ is chosen as in Proposition 2. Let $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$ be such that $\bar{h}_{i,j}=\star$, (i.e., $(i,j)$ is a non-revealed user-item cluster in $\overline{\mathbf{H}}$). Then, for all $\delta\in(0,1)$, the event: there exists $\mathcal{I}\subset\mathcal{U}^{\star},\mathcal{J}\subseteq\mathcal{V}^{\star}$ with $|\mathcal{I}|=|\mathcal{J}|=d$, such that $\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})\gtrsim m^{2}\kappa^{4},$ and
$\!\!\forall(i^{\prime},j^{\prime})\in\mathcal{I}\times\mathcal{J},\quad\bar{h}_{ij^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime}j^{\prime}}\neq\star,\ \ \bar{h}_{i^{\prime}j}\neq\star,$ (17)
@@
The proof of Proposition 3 is given in Appendix B. It is worth noting that the “non-degeneracy” property allows us to rule out pathological cases for which it may not be possible to impute the missing entries from the observed ones, at least using our approach. The need for such property is apparent in the proof, and we refer the reader to Appendix B to see how this property comes into our analysis.
4.3 The completed matrix $\widehat{\mathbf{H}}$